Juju408:

Which of the following is the correct graph of the solution to the inequality −13 > −5x + 2 > −28?

Juju408:

umm:

Step by step for: \[−13>−5x+2>−28\] First and foremost we add -2 to all parts, so it looks like this\[−13+−2>−5x+2+−2>−28+−2\]Adding it comes out to be\[−15>−5x>−30\]Now, with that, we divide all parts by -5\[\frac{ -15 }{ -5 }>\frac{ -5x }{ -5 }>\frac{ -30 }{ -5 }\]Diving that means we get \[3<x<6\]So therefore, your answer is A. (: